the rate of electron emission from 4mg of 210Pb80 with half life 5 days is

answer is 1.84×10^13


1
Expert's answer
2021-09-07T14:07:24-0400

The rate of electron emission "\dfrac{dN}{dt}" is proportional to the number of nucleus "N" in sample, accoring to the decay law (see https://en.wikipedia.org/wiki/Radioactive_decay#Universal_law):


"\left| \dfrac{dN}{dt} \right| = \lambda N"

where "\lambda = \dfrac{\ln2}{T}", where "T = 5days = 432000s" is the half life of 210Pb80.

The number of nuclei in sample with mass "m = 4mg = 0.004g" is given as follows:


"N = \dfrac{m}{M}\cdot N_A"

where "M = 210\dfrac{g}{mol}" is the molar mass of Pb, and "N_A \approx 6.022\times 10^{23}mol^{-1}" is the Avogadro number. Putting it all together, obtain:


"\left| \dfrac{dN}{dt} \right| = \dfrac{mN_A\cdot \ln2}{TM}\\
\left| \dfrac{dN}{dt} \right| = \dfrac{0.004g\cdot 6.022\times 10^{23}mol^{-1}\cdot \ln2}{432000s\cdot 210g/mol} \approx 1.84\times 10^{13}s^{-1}"

Answer. "1.84\times 10^{13}s^{-1}".